OpenStudy (anonymous):

subtract, simplify if possible 13/v-8 - 13/8-v

OpenStudy (anonymous):

\[\frac{26}{v-8}\]

OpenStudy (phi):

You need a common denominator. In this case, if you multiply the 2nd term by -1/-1 (-1 times both the top and bottom) you get -13/(v-8). Then just add (or subtract) tops: (13 - -13)/(v-8) = (13+13)/(v-8)= 26/(v-8)